OpenStudy (anonymous):

Factor the trinomial: 10u^2+17u+3

OpenStudy (anonymous):

@Luigi0210 (:

OpenStudy (luigi0210):

Start off with the basic, what multiplies to 3?

OpenStudy (anonymous):

1 and 3

OpenStudy (luigi0210):

@yakami no giving away answers

OpenStudy (anonymous):

10= 2*5 3=1*3

OpenStudy (anonymous):

Ugh ill just ignore his answer😠

OpenStudy (luigi0210):

But yea, so we got 3 and 1 (__x+3)(__x+1) Now what multiples to get 10?

OpenStudy (anonymous):

Does it matter if I write it as (_x+1)(_x+3) ?

OpenStudy (luigi0210):

Not at all

OpenStudy (anonymous):

Oh okay(: and 2 and 5 multiply to give 10.

OpenStudy (luigi0210):

And also 10*1, but of we plug that in we either get a greater value or a lesser value. So we use 5*2

OpenStudy (anonymous):

Ahhhhh I get it!!

OpenStudy (luigi0210):

Yup, so now let's plug in 2 and 5 into their correct place.

OpenStudy (anonymous):

I got 10u^2+11u+3 but its because I had wrote it as (2u+1)(5u+3)

OpenStudy (luigi0210):

So that leaves us with it being the other way around, right?

OpenStudy (anonymous):

Yes, but how do I know what numbers I need to change?

OpenStudy (luigi0210):

Factoring is mainly a trial and error process, the more practice you get the better you'll get at it. With enough practice factoring will be a breeze for you :)

OpenStudy (anonymous):

My algebra teacher said the same thing. I'm going to be working on some other problems , if I get stuck in any. Can I ask you for help?

OpenStudy (luigi0210):

Yup, of course :) but if I'm not available I'm sure @Jamierox4ev3r will be more than happy to assist you, right? ;)

OpenStudy (anonymous):

Okay then(: